OpenStudy (anonymous):

negative 5/a plus 7/a

OpenStudy (anonymous):

simplify

OpenStudy (astrophysics):

They have a common denominator so just add them as you would normally with a fraction.

OpenStudy (astrophysics):

\[-\frac{ 5 }{ a } + \frac{ 7 }{ a } = \frac{ ? }{ a }\]

OpenStudy (anonymous):

2/a?

OpenStudy (astrophysics):

Yeah!

OpenStudy (anonymous):

\[\frac{ -5+7 }{ a }\] \[\frac{ 2 }{ a }\]

OpenStudy (anonymous):

thank you lots!!

OpenStudy (astrophysics):

Np :)